OpenStudy (anonymous):

find the exact values of the cosecant and tangent of 210

OpenStudy (campbell_st):

ok... so this is a 3rd quadrant angle where tan is positive sin and cos are negative in the 3rd quadrant angles are defined by \[ang = (180 + \theta)\] then the value of theta is \[\theta = 30\] so find tan(30) and cosecant (30) = - 1/sin(30) these are both exact values.